Mean of binomial distribution is 20 and s.d is 4.find n,p,q
1
Expert's answer
2020-12-06T20:10:42-0500

Given:

mean= np=20 -------------------(1)

SD= "\sqrt{npq}" =4

that is npq= 16 -------------------(2)



Solution:


Divide equation 2 by equation 1, we get

q= "\frac{16}{np}=\frac{16}{20}=\frac{4}{5}"


and p=1-q= 1-"\frac{4}{5}=\frac{1}{5}"


n= "\frac{mean}{p}= \frac{20}{1/5}= 100"